### v1

# Maximizing domain generalization in fetal brain tissue segmentation: the role of synthetic data generation, intensity clustering and real image fine-tuning

Abstract
Fetal brain tissue segmentation in magnetic resonance imaging (MRI) is a crucial tool that supports the understanding of neurodevelopment, yet it faces challenges due to the heterogeneity of data coming from different scanners and settings, and due to data scarcity. Recent approaches based on domain randomization, like SynthSeg, have shown a great potential for single source domain generalization, by simulating images with randomized contrast and image resolution from the label maps. In this work, we investigate how to maximize the out-of-domain (OOD) generalization potential of SynthSeg-based methods in fetal brain MRI. Specifically, when studying data generation, we demonstrate that the simple Gaussian mixture models used in SynthSeg enable more robust OOD generalization than physics-informed generation methods. We also investigate how intensity clustering can help create more faithful synthetic images, and observe that it is key to achieving a non-trivial OOD generalization capability when few label classes are available. Finally, by combining for the first time SynthSeg with modern fine-tuning approaches based on weight averaging, we show that fine-tuning a model pre-trained on synthetic data on a few real image-segmentation pairs in a new domain can lead to improvements in the target domain, but also in other domains. We summarize our findings as five key recommendations that we believe can guide practitioners who would like to develop SynthSeg-based approaches in other organs or modalities.
keywords:
Fetal Brain , Synthetic Data, Domain Randomization, Segmentation, Single-Source Domain Generalization, Fine-Tuning
journal: Medical Image Analysis
[…]
booktabs
[…]
1 Introduction
1.1 Motivation
Fetal development is a crucial period of early life that plays an important role in shaping future physiological and neurological development (Halfon et al., 2014). Monitoring brain maturation during this period and detecting deviations from typical neurodevelopment is essential for the early detection of pathologies and timely intervention (Saleem, 2014). While ultrasonography (US) is the most widely used imaging modality for fetal examination due to its broad availability and low cost, fetal magnetic resonance imaging (MRI) is increasingly gaining prominence (Glenn and Barkovich, 2006). MRI is less operator-dependent than US and offers superior soft-tissue contrast, which is particularly valuable in studying fetal brain development, with numerous studies highlighting its potential in both healthy and pathologically developing fetuses (Jakab et al., 2021; Aviles Verdera et al., 2023; Machado-Rivas et al., 2022; Shen et al., 2022). Conditions such as ventriculomegaly, spina bifida, agenesis of the corpus callosum, cytotoxic and vasogenic edema, and intracranial hemorrhages can be effectively identified and assessed through MRI (Garel and Garel, 2004; Pfeifer et al., 2019). Although automated methods for fetal brain MRI analysis show particular promise in enabling rapid and reliable scan processing (Uus et al., 2023), they struggle with the same limitations as most of the approaches developed for medical imaging: limited data availability and domain shifts, which are particularly exaggerated in the fetal population (Dockès et al., 2021; Varoquaux and Cheplygina, 2022; Guan and Liu, 2022).
Figure 1: Illustration of the variations observed in fetal super-resolution reconstructed volumes causing domain shifts, presented for fetuses with different gestational ages (wGA - gestational age in weeks), MR acquisition sites and field strengths (KISPI: 1.5T, 3T, GE Healthcare; CHUV: 3T, Siemens Healthineers; KCL: 0.55T, Siemens Healthineers), and super-resolution algorithm used (IRTK, MIAL, NeSVoR) (Kuklisova-Murgasova et al., 2012; Tourbier et al., 2020; Xu et al., 2023). Both neurotypically (indicated by crosses) and pathologically (indicated by a hospital icon) developing fetuses are presented.
Domain shift, or distribution shift, occurs when machine learning (ML) or deep learning (DL) models are trained on one dataset but tested on a different one with significantly varying characteristics. In MRI, domain shifts can result from variations in the imaging protocol, differences in the scanner hardware, and manufacturer disparities (Yan et al., 2020), and in fetal MRI these parameters largely vary across centers (Kebiri et al., 2022; Xu et al., 2024; Ciceri et al., 2024). Fetal MRI poses additional challenges due to the small size of the region of interest and the high variability of brain structures, which change significantly during gestation and due to pathological processes (Stiles and Jernigan, 2010; Dubois et al., 2020) creating additional heterogeneity. Furthermore, fetal MR images are typically acquired as orthogonal stacks of two-dimensional (2D) slices to mitigate motion artifacts, necessitating super-resolution (SR) reconstruction algorithms to create a high-resolution three-dimensional (3D) volume (Rousseau et al., 2010; Gholipour et al., 2010; Kuklisova-Murgasova et al., 2012; Tourbier et al., 2015; Ebner et al., 2020). SR algorithms introduce further variability through reconstruction artifacts, use of skull stripping, and differences in image appearance and contrast, adding another layer of domain shift (Xu et al., 2023). These compounded shifts complicate training models that perform well and consistently across different datasets. Figure 1 illustrates the variability of fetal MR images related to changes in the brain morphology across maturation, acquisition parameters, magnetic field strength, and SR methodology used.
A recent study by Payette et al. (2024a) presenting the Fetal Tissue Annotation and Segmentation (FeTA) 2022 MICCAI challenge results shows that the aforementioned domain shifts can degrade the performance of the models when they are trained and tested on different clinical datasets. Their results reveal drops in performance occurring for white matter (WM), grey matter (GM) and ventricles segmentation in the out-of-domain (OOD) testing compared to in-domain testing. This emphasizes the importance of tackling domain shifts in fetal MRI analysis, induced by high morphological variability of the fetal brain, the MRI instrumentation and acquisition, and SR algorithms.
Many techniques have been developed to address domain shifts, the most widely used being data augmentation (Shorten and Khoshgoftaar, 2019), which aims at altering the images in a controlled manner before learning to make the model more robust to heterogeneous data. Recently, however, domain randomization (Tobin et al., 2017) techniques have seen great success in the field of MR image analysis (Billot et al., 2021, 2023b; Liu et al., 2024a). These methods begin with label maps rather than real images, generating synthetic images with randomized contrasts using Gaussian mixture models, alongside common perturbations such as geometric deformations and resampling that simulate different image resolution and partial volume effects. This strategy has demonstrated excellent OOD generalization performance, especially in MRI, a modality often challenged by substantial contrast variability across centers, scanners, and sequence variants (Billot et al., 2023b).
1.2 Contributions
In this work, we aim to investigate how to make the best use of the synthetic-data-driven domain randomization approaches (Tobin et al., 2017) for fetal brain MRI tissue segmentation. We study how various sources of synthetic and real data impact the model’s performance in a single-source domain generalization (SSDG) setting. In this setting, training data come from only one domain (source), and the objective is to generalize to any potential testing (target) domain.
This paper presents a comprehensive examination of domain randomization methods in fetal brain MRI, with a particular focus on SynthSeg-based approaches. Our contributions can be summarized as follows:
(i) We train multiple variants of SynthSeg across datasets from different acquisition sites, with varying field strengths, processing methods, and image quality. We investigate how two key components of our fetal-tailored model, FetalSynthSeg (Zalevskyi et al., 2024)—namely intensity clustering and the use of meta-classes (where tissues with similar intensities are grouped as one label)—improve OOD generalization.
(ii) We explore the use of physics-based intensity generation methods, such as numerical phantoms like FaBiAN (Lajous et al., 2022), as an alternative to Gaussian mixture models for intensity sampling. This allows us to study the effectiveness of achieving contrast invariance using more physically-grounded simulations.
(iii) We show how data quality significantly impacts the generalization ability of synthetic methods, emphasizing the need to carefully consider the quality of training data in domain generalization tasks.
(iv) Finally, for the first time, we combine SynthSeg with modern fine-tuning approaches based on weight-space averaging (Rame et al., 2022; Wortsman et al., 2022b; Gagnon-Audet et al., 2023). We demonstrate that fine-tuning SynthSeg on even a small number of real images from a new domain can improve generalization both in-domain and out-of-domain, providing a flexible method to boost performance across varying datasets.
The code, the trained model, and a Docker image with the model are available at https://github.com/Medical-Image-Analysis-Laboratory/FetalSynthSeg.
2 Background
2.1 Tackling domain shifts
Over the years, multiple approaches that follow different strategies have been proposed to tackle domain shifts. For instance, Karani et al. (2018) introduced a transfer learning method involving the learning of batch normalization parameters for each domain while sharing convolutional filters across domains. This allowed models to adapt to new domains by fine-tuning specific parameters without altering the core network architecture. Similarly, Zhou et al. (2021) proposed a style transfer method that addressed domain shifts by mixing styles of source domains to implicitly synthesize novel domains, thereby diversifying training data. Meta-learning has also shown promise in domain generalization. Li et al. (2018) developed a model-agnostic meta-learning framework where source and target domains were simulated during training by selecting random subsamples from available source domains. This enabled the model to learn to generalize across different domains by optimizing for learning of domain-invariant features. Additionally, Hu et al. (2023) explored data harmonization techniques to reduce domain shifts by standardizing data across different sources, thereby mitigating variability that could lead to performance degradation.
While these methods have been effective in various medical imaging tasks, they often necessitate a dataset of target domain images and, in some cases, annotations. In fetal MRI, obtaining such datasets can be costly and challenging  (Payette et al., 2023). The manual segmentation of a single case often takes several hours of expert work (e.g., eight hours per case in Kyriakopoulou et al. (2017)). Furthermore, methods that rely on multiple source domains to learn domain-invariant features are limited when only a single training domain is available or when the diversity of available source domains does not adequately represent the target domain  (Zhou et al., 2023). This limitation is especially pronounced in fetal imaging, where newly emerging scanner field strengths or SR algorithms can introduce previously unseen domains. The poor performance in these new unseen domains creates obstacles in adopting low-field ($<$1T) and ultra-low-field ($<$0.1T) MR systems, which hold great potential for democratizing access to advanced neuroimaging for broader populations (Aviles Verdera et al., 2023). The scarcity and small size of existing fetal brain imaging datasets further exacerbate this issue (Payette et al., 2023; Lajous et al., 2022).
2.2 Single-source domain generalization (SSDG)
Given these challenges, SSDG techniques are particularly relevant in the medical imaging domain. These techniques aim to train models on a single source domain while ensuring they can generalize well to unseen target domains. Common approaches used in medical image analysis include altering image appearance through texture and intensity augmentation and blending generated views. Ouyang et al. (2022) used randomly weighted shallow convolutional neural networks in a global intensity non-linear augmentation (GIN) module to augment images used to learn a segmentation task. Similarly, Li et al. (2023) simulated frequency discrepancies between domains to achieve SSDG by mixing different frequency views of simulated images and incorporating a self-supervised task to learn context-aware representations.
Other approaches like the adversarial domain synthesizer (ADS) by Xu et al. (2022) synthesized new domains aimed to cover a large area of plausible distributions. They introduced a mutual information regularizer to enforce semantic consistency between synthetic domain images. Liu et al. (2022) used dictionary learning to extract prior information on the semantic shape of segmentations and images, utilizing this during test-time augmentation to improve model generalizability.
Despite advancements, SSDG still faces challenges due to the lack of diversity in training data. Current methods are limited by the semantic information in existing data and primarily focus on 2D data, with extensions to 3D image processing still unexplored. At the same time, the FeTA Challenge 2021 demonstrated that using 3D models for fetal brain tissue segmentation results in better performance than 2D models, highlighting the need for SSDG approaches applicable to 3D images (Payette et al., 2023).
2.3 Synthetic data and domain randomization
Synthetic data generation and domain randomization have emerged as promising solutions to SSDG limitations  (Al Khalil et al., 2023; Paproki et al., 2024; Gopinath et al., 2024).
Synthetic data can be generated using deep learning approaches like generative adversarial networks (GANs) or diffusion models, though these methods are challenging to train and require extensive datasets to learn the latent representation of the data (Kazerouni et al., 2023; Kazeminia et al., 2020). Alternatively, numerical phantoms like FaBiAN have been developed to generate images based on physical simulations and anatomical priors such as tissue segmentation, without extensive pre-training (Lajous et al., 2022). FaBiAN, the first Fetal Brain MR Acquisition Numerical phantom, can generate multiple MR images for the same subject with varying acquisition parameters and model stochastic fetal movements, producing highly realistic images that allow to complement scarce clinical datasets (Bhattacharya et al., 2024; Lajous et al., 2024b, a).
In contrast, SynthSeg, proposed by Billot et al. (2021), avoids physical simulation altogether and generates a highly diverse synthetic dataset of images achieving domain randomization. It uses anatomical priors but simulates tissue intensities fully randomly, creating images with appearances that often exceed realistic. This strategy assumes that the diversity of such randomly generated data will cover the constrained diversity existing between real domains. SynthSeg has been widely used for brain MRI analysis across various tasks, including stroke segmentation, infant and adult skull stripping, infant brain tissue segmentation and contrast-agnostic extraction of anatomical representations of healthy and pathologically developing adult brains (Omidi et al., 2024; Kelley et al., 2024; Liu et al., 2024b, a; Shang et al., 2022; Valabregue et al., 2024; Gopinath et al., 2024).
2.4 Fetal brain tissue segmentation
While many synthetic data-driven approaches discussed above primarily focus on adults or infants, this work addresses fetal brain tissue segmentation, which presents unique challenges related to in utero imaging, including limited tissue contrast compared to postnatal imaging, and often fewer annotation classes than in adult MRI complicating the simulation of the heterogeneous fetal brain tissues.
The FeTA Challenge, organized as a satellite event of the MICCAI conference, plays a crucial role in advancing the field by offering a platform for evaluating and benchmarking segmentation algorithms tailored for fetal brain imaging (Payette et al., 2024a). The 2022 challenge highlighted the importance of data augmentation for generalizing across diverse multi-site datasets. The top-performing methods used 3D DL architectures, with nnU-Net being the most popular. The best-performing team achieved a global average Dice score (mDSC) of 0.816±0.11 and a 95th percentile Hausdorff distance (HD95) of 2.347±2.51 mm in OOD evaluations.
Using FeTA’s publicly available data, Huang et al. (2023) proposed a hybrid convolution-transformer network, achieving a Dice score (DSC) of 0.837±0.03 on a held-out in-domain validation set. Another notable approach, BOUNTI, combines 3D U-Net and Attention U-Net architectures and was trained on a diverse 380 multi-site fetal brain MR images from the developing Human Connectome Project, reaching a DSC of 0.89±0.02 (Uus et al., 2023). However, BOUNTI’s private training dataset with higher quality annotations and distinct annotation scheme limits the direct comparison with models trained on FeTA data.
These results underscore the need for further research into in utero fetal brain segmentation and the potential of synthetic data to bridge the domain shifts appearing when models are trained on limited single-domain data. Our work aims to address these gaps.
3 Materials and methods
In this section, we describe the data and approaches used for learning fetal brain tissue segmentation in our experiments.
3.1 Data
This study utilizes open-access data from the FeTA challenge as well as private clinical datasets from two institutions, as summarized in Table 3.1. All acquisitions were performed without the use of maternal or fetal sedation. In all datasets, several volumes were acquired for each subject using a 2D single-shot fast-spin-echo sequence in at least three orthogonal orientations (axial, coronal, and sagittal), to mitigate fetal motion artifacts (see Acquisition Parameters column in the Table 3.1). A super-resolution reconstruction algorithm was used afterward to merge them into a single high-resolution volume suitable for 3D volumetric segmentation.
Table 1: Dataset properties. $N_{n}$ number of neurotypical subjects, $N_{p}$ number of pathological subjects. KISPI* - publicly available dataset of the FeTA challenge (Payette et al., 2021)
{tblr}
Q[c,m]Q[c,m]Q[c,m]Q[c,m]Q[c,m]Q[c,m]Q[c,m] [1.5pt] Site Scanner Acquisition
Parameters
SR
algorithm
SR
resolution
($mm^{3}$)
GA
$N_{n}/N_{p}$
[1pt] KISPI* GE Signa
Discovery
MR450/MR750
(1.5T/3T) SS-FSE
TR/TE:
2500-3500/120 ms
0.5 x 0.5 x 3.5 $mm^{3}$ MIAL
IRTK $0.5^{3}$
$0.5^{3}$ 20-34
20-34 25/15
24/16
CHUV Siemens
MAGNETOM
Aera
(1.5T) HASTE
TR/TE:
1200/90 ms
1.13 x 1.13 x 3 $mm^{3}$ MIAL $1.1^{3}$ 21-35 25/15
KCL Siemens
MAGNETOM
FREE.MAX
(0.55T) HASTE
TR/TE:
2600/106 ms
1.45 x 1.45 x 4.5 $mm^{3}$ SVRTK
NeSVoR $0.8^{3}$ 21-35 40/0
[1.5pt]
FeTA Challenge Dataset. This dataset includes images of 80 subjects scanned as a part of clinical acquisitions at the University Children’s Hospital Zurich (KISPI) between 2016 and 2019 (Payette et al., 2021; Payette and Jakab, 2021b). T2-weighted single-shot fast spin echo (SS-FSE) sequences were acquired on a 1.5-T or 3-T GE whole-body scanner, either using an 8-channel cardiac coil or body coil. The acquired images have an in-plane resolution of $0.5\times 0.5mm^{2}$ and a slice thickness of 3-5 mm. Both healthy and pathologically developing subjects were imaged. Half of the dataset was reconstructed with MIALSRTK and the other half with simple-IRTK SR algorithms, creating high-resolution volumes with an isotropic resolution of $0.5mm^{3}$ (Tourbier et al., 2020; Kuklisova-Murgasova et al., 2012). The ethical committee of the Canton of Zurich, Switzerland approved the prospective and retrospective studies that collected and analyzed the MRI data presented in this dataset (Decision numbers: 2017-00885, 2016-01019, 2017-00167), and a waiver for ethical approval was acquired for the release of a fully anonymous dataset for research purposes (Payette et al., 2023). Due to reported quality issues with a part of the dataset (Payette et al., 2021), manual quality control of the reconstructed volumes was performed by MBC, an engineer with 20 years of experience in neuroimaging, following the guidelines and tools from Sanchez et al. (2024). A global quality ranking (continuous rating between 0 and 4) was reported, with images rated lower than 1 deemed to have very poor quality (see Figure 2(Payette et al., 2021).
Figure 2: Distribution of quality control scores for the KIPSI-irtk and KISPI-mial datasets. The dotted red line at 1 represents a classification threshold below which SR reconstructions are assessed of ’poor’ quality.
CHUV Clinical Dataset. Acquired at Lausanne University Hospital (CHUV) between 2013 and 2021, using a 1.5 T MR scanner (MAGNETOM Aera, Siemens Healthcare), this dataset is composed of 40 subjects imaged with a T2-weighted Half-Fourier Acquisition Single-shot Turbo spin Echo (HASTE) sequence, with an 18-channel body coil and a 32-channel spine coil. Original images have a voxel size of 1.13 × 1.13 × 3.00 $mm^{3}$, an inter-slice gap of 10%, and an acquisition time between 26 and 36 seconds. They were further reconstructed with MIALSRTK (Tourbier et al., 2020) to a 3D isotropic $1.1mm^{3}$ volume. All images were anonymized and the dataset collection was approved by the ethics committee of the Canton de Vaud (CER-VD 2021-00124).
KCL Low-Field Clinical Dataset. Collected with a 0.55-T scanner (MAGNETOM FREE.MAX, Siemens Healthineers) at St Thomas’ Hospital, London, this dataset includes images acquired using a HASTE sequence with the following parameters: TR/TE of 2500ms/106ms and a voxel size of 1.45 × 1.45 × 4.5mm³, for a total acquisition time between 64 and 122 seconds. Acquired volumes, if deemed of sufficient quality, were used for SR reconstruction with both NeSVoR and SVRTK algorithms of a single high-resolution isotropic volume of $0.8mm^{3}$ (Xu et al., 2023; Uus et al., 2022). The data, acquired between 2019 and 2021, were anonymized and approved for sharing by the Ethics Committee Bromley (21/LO/0742). No ground truth segmentation labels are available for this dataset, therefore it was not used for synthetic data generation or segmentation learning.
Pre-processing and labeling. All SR reconstructed volumes were zero-padded to the same image size (256x256x256 voxels) and were labeled using the FeTA annotation scheme when applicable (Payette et al., 2021). Each case includes an annotated label map with seven brain tissue types: WM, GM, cerebrospinal fluid (CSF), ventricles, deep GM, cerebellum, and brainstem. During the model training, all input images and segmentation were further resampled to 0.5 x 0.5 x 0.5 mm³ and cropped or padded to the same size ($256^{3}$ voxels).
3.2 Methods
Our experiments will rely on three components: SynthSeg (Billot et al., 2021), the numerical phantom FaBiAN (Lajous et al., 2022), and FetalSynthSeg, our tailored version of SynthSeg for fetal brain MRI segmentation (Zalevskyi et al., 2024). We introduce each of these elements in more detail in the following.
3.2.1 SynthSeg
SynthSeg is a domain randomization approach that leverages anatomical segmentations rather than real images to produce synthetic data. Given a dataset of segmentations $\{S_{n}\}$, it randomly samples a segmentation from this dataset which serves as an initial anatomical prior, applies spatial augmentations $\phi$, and generates synthetic images $G$ at a location $(x,y,z)$ by sampling a Gaussian Mixture Model (GMM) defined for a label $L$ with fully randomized parameters:
$G(x,y,z)\sim\mathcal{N}(\mu_{L(x,y,z)},\sigma_{L(x,y,z)}^{2}),$
where $\mu_{L}$ and $\sigma_{L}$ are sampled from uniform distributions $\mathcal{U}(a_{\mu},b_{\mu})$ and $\mathcal{U}(a_{\sigma},b_{\sigma})$, defined by upper and lower bounds $a,b$, respectively. The image is then deformed using both affine and non-rigid diffeomorphic transformations and corrupted by adding a bias field, intensity transformations and by simulating various image resolutions. By corrupting the images much more significantly than what would be realistic, SynthSeg ensures that the learned features will be robust to strong corruption and shifts occurring in the heterogeneous data found in practice.
Implementation. We utilize the Pytorch version of the generator and set all of its hyperparameters to values from the Brain-ID implementation, another domain randomization approach based on SynthSeg that uses synthetic data to learn robust brain representations (Liu et al., 2024a).
3.2.2 FetalSynthSeg
Fetal brain MR images exhibit a high degree of heterogeneity and typically have a small number of segmentation classes available (e.g., seven in FeTA). Solely relying on a few annotation classes leads to images that miss finer structures, either within a class or in the background. Billot et al. (2021) noted this issue in cardiac segmentation and proposed to use intensity clustering on the background of the original (intensity) image, as well as within the different classes. This allows to simulate finer variations in the data. In FetalSynthSeg (Zalevskyi et al., 2024), we took this approach one step further by defining four meta-classes containing tissues with similar intensities in the original images. The meta-classes are: white matter (white matter, cerebellum, and brainstem), gray matter (cortical and deep gray matter), cerebrospinal fluid (ventricles and external CSF), and non-brain tissues (skull, uterus, fetal body, and surrounding maternal tissue) which include all non-zero image voxels that do not belong to any ground truth segmentation label. Assigning a meta-class to non-brain tissues is essential to effectively simulate the background environment surrounding the fetal brain and account for variations in SR algorithms’ performance and quality in skull stripping, ensuring that the model is trained on data that reflect realistic preprocessing variability common in fetal MRI. During the generation process, after randomly selecting the subject that will be used for generation, we first aggregate its labels into these four meta-classes. Each meta-class is then divided into a random number of subclasses, sampled from a uniform distribution $\mathcal{U}(a_{subcl},b_{subcl})$. The splitting into subclasses is performed using the Expectation Maximization (EM) algorithm, which is conditioned on the original intensity image (Dempster et al., 1977). This process is illustrated in Figure 3.
Using meta-labels and splitting them into subclasses, rather than relying on the original segmentation labels, helps prevent artificial boundaries from appearing in synthetic images that align with segmentation boundaries. This approach avoids the risk of the model learning shortcuts by relying on distinct label contrasts that may not actually exist in real acquisitions (for example as with white matter and cerebellum that have similar intensity distributions in fetal brains (Payette et al., 2024b)).
Figure 3: FetalSynthSeg pipeline for meta-label splitting into sub-classes. First original segmentation labels are merged into four meta-labels (WM, GM, CSF, non-brain tissue). Then, each meta-label is randomly split into sub-classes. Finally, an independent GMM is sampled for each of the sub-classes and used to sample intensities for voxels inside of it.
Implementation. We use the same generator hyperparameters as in the SynthSeg experiments described in Section 3.2.1, and the same implementation based on Brain-ID (Liu et al., 2024a), while enhancing it with the discussed meta-classes.
3.2.3 FaBiAN
The Fetal Brain magnetic resonance Acquisition Numerical phantom (FaBiAN) is an open-source tool designed to simulate clinical T2-weighted fast spin echo (FSE) sequences of the fetal brain (Lajous et al., 2022). Like FetalSynthSeg, it is based on brain tissue segmentation as an anatomical prior of the fetal brain to simulate and merge the segmentation labels into coarse tissue classes (WM, GM, and CSF). However, instead of using simple GMMs to model intensity distributions of voxels belonging to these different classes, FaBiAN employs the extended phase graph (EPG) formalism to simulate the signal formation process in FSE acquisitions. As such, it estimates the T2 decay over time in every voxel of the fetal brain based on the T1 and T2 relaxation times known from the literature of each of the three tissue classes, while accounting for sequence-specific mechanisms. We have modified the code of FaBiAN to: i) include the non-brain tissue meta-class simulated with the same subclasses as FetalSynthSeg, and ii) use randomized T1 and T2 values coming from a reference distribution based on the real values reported in the literature, which are then integrated into the physical simulation.
We also consider a randomized version of FaBiAN (randFaBiAN), where instead of using reference T1 and T2 values for the physical simulation, all values are sampled from a broad, unrestricted distribution. This approach randomizes the physical properties of the brain tissues, making it more akin to FetalSynthSeg’s method.
Table S5 in the supplementary material details the parameters used in FaBiAN and randFaBiAN, and Figure 4 offers a visual comparison of the images generated by the discussed algorithms.
Implementation. We use a Matlab implementation of FaBiAN_v2 (Lajous et al., 2024a, b) wrapped in a Python interface available on our GitHub page github.com/Medical-Image-Analysis-Laboratory/fabian_utils.
Figure 4: Examples of the images generated by: (B) FaBiAN, (C) randFaBiAN, and (D) FetalSynthSeg based on the same (A) original image and its segmentation.
3.3 Training on real images
To assess the domain generalization capabilities of the models, we compare them to a baseline model trained exclusively on real images from the same training split. For this baseline model, we use the exact same set of real images corresponding to the segmentations used to generate synthetic images in all the other generative approaches. To ensure a fair comparison, we apply the same augmentation pipeline during training as was used for the models trained on synthetic data (including random resampling, spatial deformations and artifact generation). This setup allows us to isolate the impact of randomized intensity and contrast generation, as all other transformations—apart from intensity sampling—are applied with identical parameters across both the baseline and synthetic-data-driven models.
3.4 Fine-tuning and weight interpolation
Finally, we explore how integrating synthetic and real data can improve the performance and robustness of segmentation models. Previous studies have shown that synthetic data is particularly valuable when real data is scarce, often achieving comparable or superior performance  (Khosravi et al., 2023; Thambawita et al., 2021). Additionally, Frid-Adar et al. (2018); Loizillon et al. (2023) demonstrated that pre-training on synthetic data followed by fine-tuning with real data leverages the strengths of both data types. However, with limited real training data, fine-tuning can lead to overfitting and degrade out-of-distribution performance  (Kumar et al., 2022). To address this, we explore weight-space ensembles  (Rame et al., 2022; Wortsman et al., 2022a; Gagnon-Audet et al., 2023), where models pre-trained on synthetic data ($\theta_{\text{synthetic}}$) are fine-tuned with real data ($\theta_{\text{real}}$), and their weights are linearly interpolated during inference with an interpolation rate $\alpha$ :
$\theta_{\text{interpolated}}=(1-\alpha)\theta_{\text{synthetic}}+\alpha\theta_{\text{real}}$
This approach has been shown to enhance out-of-distribution robustness while maintaining or improving in-distribution accuracy, which is particularly important in our experimental set-up  (Wortsman et al., 2022b; Stojanovski et al., 2022).
3.5 Experimental setting
Experiment layout. Our experiments are structured in three sections. (i – Cross-validation) First, we evaluate the out-of-domain generalization of FaBiAN, randFaBiAN, SynthSeg, FetalSynthSeg and real training data in fetal brain MRI in a cross-validation setting, using the FeTA (KISPI-IRTK, KISPI-MIAL) and CHUV-MIAL datasets. (ii – Pure testing) Secondly, we conduct a further evaluation on the held-out 0.55T test data from the KCL, reconstructed using SR reconstruction methods not seen during training. As we do not have labels for this data, we construct growth charts of healthy subjects and evaluate the deviation of the estimated volumes of tissues from the reference volumes obtained from the manual segmentations of healthy subjects from the FeTA dataset. (iii – Fine-tuning) Finally, as fine-tuning techniques are a common tool to adapt deep learning techniques to new domains, we investigate how fine-tuning and weight-space averaging (Rame et al., 2022; Wortsman et al., 2022b; Gagnon-Audet et al., 2023) can be leveraged to improve the performance of synthetically trained models on new domains.
Models architecture and training. In this study, we primarily aim to investigate how we can best use synthetic data generation methods for fetal brain MRI tissue segmentation. For all evaluations, we employ a 3D U-Net model, similar to the architectures in Billot et al. (2021) and Valabregue et al. (2024) implemented in Pytorch with MONAI (Paszke et al., 2019; Cardoso et al., 2022). The model includes five levels with 32 feature maps at the first level, doubling with each layer, and utilizes 3x3x3 convolutions with LeakyReLU activations, except the final layer which uses softmax. Skip connections are included. We chose to focus on using a simple model and evaluating multiple data generation approaches rather than trying to build the most competitive version of synthetic data generation using larger models, ensembling or post-processing.
Training is done using the Adam optimizer with a learning rate of $10^{-3}$ and a loss function that combines Dice and Cross-Entropy losses, as used by Valabregue et al. (2024). Additionally, we employ the ReduceLROnPlateau scheduler (with a factor of 0.1 and a patience of 10 epochs) and EarlyStopping (with patience of 100 iterations). All experiments are conducted with a batch size of 1, implemented using the PyTorch Lightning (Paszke et al., 2019) framework, and executed using Nvidia RTX 3090 GPUs.
For SynthSeg and FetalSynthSeg, we generate synthetic images on the fly, training for up to 80,000 iterations. For FaBiAN and randFaBiAN, we created an offline database of 50 synthetic images per real image, resulting in 6000 synthetic images in total, and trained for a maximum of 80,000 iterations. Due to computational constraints, on-the-fly generation of FaBiAN images was not feasible since the simulator, implemented in Matlab, required an average of 287 seconds to generate a single synthetic image compared to 1 second needed to generate a synthetic image with SynthSeg or FetalSynthSeg. For experiments involving real images, the training was performed for a maximum of 500 epochs to avoid overfitting.
Data augmentation. We use the data augmentations from SynthSeg for all of the models trained. They contain random non-linear deformation, random contrast change (Gamma transformation), random Gaussian noise, random bias field, and a random resampling simulating an acquisition at a different resolution.
Additional baselines. We also include, when applicable, comparisons to fit_nnUnet, a state-of-the-art approach that won the FeTA 2022 challenge  [^3] The model is available on the FeTA challenge DockerHub page Payette et al. (2024b).. It is based on an ensemble of nnUnet models trained with different augmentations, and denoising auto-encoders as a postprocessing step to correct poor segmented cases. It was trained using the full FeTA challenge dataset, which included 120 images including KISPI-irtk and KISPI-mial datasets while also an additional 40 subjects from another institution. In our evaluation, it is considered as an upper bound of performance, achieved by the state-of-the-art algorithms trained on publicly available data.
Fine-tuning. When performing fine-tuning, we use a pre-trained FetalSynthSeg model and fine-tune it on a random subset of real images (1, 3 or 5) for a maximum of 300 iterations using a small learning rate of $10^{-6}$ to avoid overfitting. We apply the same set of data augmentations as when training on real images. Finally, we perform weight interpolation between the fine-tuned model and the pre-trained FetalSynthSeg model and explore the optimal interpolation rate $\alpha$ that achieves the best out-of-domain generalization.
Table 2: Mean DSC for models trained on different sources of data. The variance is computed over all testing subjects within a split. The best-performing method is shown in boldface in each column, and the second best is underlined. Stars are added to the best method if it was found to be statistically significantly different (Wilcoxon rank-sum test, $p<0.05$) than the second best method in the same column.
{tblr}
[…]
randFaBiAN 79.1±2.3 78.2±2.9 55.1±12.7 68.9±7.8 60.7±7.3 68.1±14.7 68.3±13.8
SynthSeg 75.9±3.9 73.7±3.9 70.9±9.2 74.8±7.8 60.5±15.7 63.4±16.8 72.2±13.0
[…]
Figure 5: Ablation study on the number of subclasses used. For SynthSeg, the original 7 labels used as generation classes are randomly split into 1-9 subclasses, while in FetalSynthSeg the meta-labels are split. We report mean DSC across all tissues for each combination of training-testing splits. The dashed line gives the average trend.
Cross-validation evaluation. We compare the performance of all models in a single-source domain generalization scenario. For this, we perform cross-validation on the labeled datasets (KISPI-MIAL, KISPI-IRTK and CHUV-MIAL), where a single dataset is used for training, and the two others for testing. Each of the training domains contains 40 images, and we randomly assigned 35 images (and all synthetic data derived from them) to training and used 5 for internal validation. All 40 images from other domains were used for testing, and the results were both reported domain-wise as well as globally aggregated.
Metrics. We report mean Dice scores across all seven segmented tissues in the paper and provide detailed results with per-tissue Dice scores and 95th percentile Hausdorff distance in the Supplementary Materials Tables S3 and S4.
Statistical analysis. We use a non-parametric Wilcoxon rank-sum test with Bonferroni correction to statistically analyze the results and compare the performance of the same subjects using different methods. A p-value for statistical significance is set to 0.05.
Figure 6: Performance comparison between FetalSynthSeg and SynthSeg (A). The aggregated performance of both models on the testing splits depending on the status of the evaluated subject (healthy/pathological) (B). Effect the quality of an image has on the performance of the models on the KISPI-mial split (C).
Figure 7: Segmentation results for models trained on the KISPI-MIAL data split, evaluated on a subject from the KISPI-IRTK split (top; excellent quality) and on another from the CHUV-mial split (bottom; acceptable quality). Arrows point to regions of substantial discrepancies across models.
4 Results
4.1 Cross-validation experiment
The global comparison of the methods is available in Table 3.5. Solely considering the performance of the methods, on average FetalSynthSeg outperforms all competitors by a large margin, showing how synthetic data generation can help bridge the gap between domains. Furthermore, when comparing FetalSynthSeg with fit_nnUNet, the winner of FeTA challenge 2022 (Payette et al., 2024a), on the CHUV-mial split (as the fit_nnUNet model was trained using both other splits), we see a competitive result of FetalSynthSeg, with a difference of only 0.6 Dice points, even if FetalSynthSeg was trained on 3 times fewer data.
Several more insights drawn from these results are presented in the following.
4.1.1 Intensity clustering is key to achieving a strong performance
As the FeTA segmentations contain only seven classes, using just the initial segmentation does not allow to fully capture the intricate details of real images. Using the original image and generating additional classes through intensity clustering allows us to tackle this issue. This is what we show in Figure 5. This figure shows how performance varies for SynthSeg and FetalSynthSeg depending on the number of sub-classes used for intensity clustering. SynthSeg with one sub-classes is the default model. There is a clear benefit of using multiple sub-classes per tissue, independently of the choice of using meta-labels. Nevertheless, we see that the meta-labels allow FetalSynthSeg to retain a small edge over SynthSeg. A statistical analysis reveals that FetalSynthSeg still outperforms SynthSeg in 7 out of 12 training splits, with SynthSeg performing better in 3 splits, and no significant difference in performance found in the remaining 2 splits.
4.1.2 Data quality greatly impacts generalization performance
Table 3.5 also shows a large variance in performance across splits, but also across subjects, with a variance up to 16.0 dice points for the CHUV-mial trained FetalSynthSeg model tested on KISPI-mial data. Focusing on SynthSeg and FetalSynthSeg, Figure 7 shows the heterogeneity from the different testing splits. It is very clear, especially for KISPI-MIAL, that pathological subjects can be much harder to successfully segment than healthy ones (Fig 7(B)), which is mostly correlated to image data quality: all methods fail at segmenting reconstructed images that had been rated to have poor quality. This is clearly illustration by Figure 7 However, data quality does not only impact the evaluation: models trained using the lower quality KISPI-MIAL data also lead to a lower generalization performance across all splits (cf. Table 3.5).
Figure 8: Tissue volumes across GA for KCL images reconstructed with SVRTK (top row) and NeSVoR (bottom row). Volumes based on manual segmentations from the FeTA dataset are shown in black, those estimated from FetalSynthSeg predictions are in green, and those estimated from segmentations produced by a model trained on real data are in blue. All models were trained on the same images or their synthetic derivatives from the KISPI-IRTK split and evaluated on 40 healthy subjects from the KCL dataset. The lines represent second-order polynomial fits, and the reported mean-squared error (MSE) is calculated between the polynomial fits of the reference (FeTA) and each model.
Figure 9: Segmentation results for models trained on the KISPI-IRTK data split, evaluated on two subjects from KCL (top; reconstructed using NeSVoR – bottom; reconstructed using SVRTK) and another from the CHUV-mial split (bottom; acceptable quality). Arrows point to regions of substantial discrepancies across models.
Figure 10: Mean DSC for models trained on the source domain. FetalSynthSeg – trained only on synthetic data from the source domain. FT on Source – are FetalSynthSeg models trained first on the synthetic data and then fine-tuned on the same source domain. Interpolated – weight averaging of the two first models.
Figure 11: Improvement in DSC of the fine-tuned target domain data compared to the SynthSeg performance. We evaluate how fine-tuning with different numbers of images {1, 3, 5} affects the performance of the fine-tuned model as well as its interpolated version with different interpolation weights. We evaluate the performance both in the target domain and in other domains. Fine-tuning weight $\alpha$ is the weight that is used for averaging: $\theta_{\texttt{interpolated}}=(1-\alpha)\theta_{\texttt{FetalSynthSeg}}+\alpha\theta_{\texttt{fine-tuned}}$. The original FetalSynthSeg is at $\alpha=0$ and vanilla fine-tuning is at $\alpha=1$.
4.1.3 SynthSeg-based approaches outperform physics-based simulation for model training
Table 3.5 also features an ablation study on four different variants for synthetic generation. The results show that FaBiAN, which restricts intensity simulation to reference tissue relaxometry properties generalizes worst (63.3 DSC). randFaBiAN simulates a wider range of T1 and T2 tissue values, and largely increases its performance to (+5; 68.3 DSC). SynthSeg then moves from using FaBiAN as a simulator to GMM gains and allows to get an additional edge (+3.9; 72.2 DSC). The final gain comes from using meta-classes (+2.7; 74.9 DSC). The role of SynthSeg-based data augmentations is also investigated in the Supplementary Experiment S1, where we show how it improves the performance of models even when using real images and is superior to the standard augmentations based on MONAI that are commonly used in the field.
These results suggest that there is a gap between FaBiAN-based approaches and FetalSynthSeg. This is further supported by Supplementary Experiment S2, where a FetalSynthSeg model trained using $6000$ samples like FaBiAN still outperforms it by 3.9 Dice points. It is also important to remember that FaBiAN-based contrast simulation requires very heavy and slow computations (around 280 seconds per subject), compared to the simple and extremely efficient contrast simulation using GMMs. Expensive physics-based simulators like  making it impractical for the online generation used in SynthSeg-based approaches.
Qualitative analysis in Figure 7 illustrate these findings, showing that models with randomized contrast generators like randFaBiAN and FetalSynthSeg are more robust to domain shifts, particularly when there is a mismatch in skull-stripping quality between training and testing data.
4.2 Pure testing experiment
4.2.1 FetalSynthSeg allows robust out of domain generalization
Figure 8 shows growth charts comparing FetalSynthSeg to a model trained on real data. The data were acquired on a 0.55T scanner at KCL, and reconstructed with two unseen SR algorithms (SVRTK (Kuklisova-Murgasova et al., 2012; Uus et al., 2022) and NeSVoR (Xu et al., 2023)). Although a new method, SVRTK is conceptually close to IRTK, and we expect KCL-SVRTK performance to show less performance drop than KCL-NeSVoR. The latter represents a double domain shift: scanner-field-strength and SR algorithm. The results in Figure 8 show that the volume estimations from FetalSynthSeg for different tissues, as well as the whole cerebral brain volume, are very close to the values estimated based on the ground truth (FeTA) for both KCL-SVRTK and KCL-NeSVoR datasets. On the other hand, the volume estimates for the model trained with real data tested on the KCL-NeSVoR dataset deviate very largely from the ground truth, highlighting the effect of the domain shift that degrades the model’s performance. This is quantified by computing the mean-squared error between the fitted curves. For KCL-SVRTK, the MSE for the model trained on real data is $639.6$ and FetalSynthSeg is $415.1$. For KCL-NeSVoR, the model trained on real data shows an averaged MSE of $12274.1$ compared to $1781.3$ for FetalSynthSeg. The difference is particularly marked for the estimation of brain, white matter and CSF volumes. Qualitative comparisons shown in Figure 9 highlight also the challenge posed by NeSVoR reconstructed volumes, clearly visible for the model trained on real data, where larger topological inconsistencies were seen, especially on the segmentation of CSF.
4.3 Fine-tuning experiment
Our last series of experiments are investigating if combining real and synthetic data could further help the out-of-domain generalization of the models. We first investigate whether using the data from the training (source) domain itself could further improve the generalization performance of the model. Then, we investigate how using a few data points from the target domain helps generalization. Throughout the experiments, we compare standard fine-tuning and weight averaging techniques.
4.3.1 Source domain fine-tuning marginally improves generalization performance of FetalSynthSeg
Figure 10 shows the results obtained when fine-tuning FetalSynthSeg on real data from the training domain and compares it to the weight-space interpolation. The results show that fine-tuning with real images from the same domain that was used for pre-training with synthetic data often does not significantly improve generalization performance and that it can even lead to performance drops in some cases. These limitations are mostly overcome by averaging the weights of the original and fine-tuned models. Overall, the weight-interpolated models show a statistically significant improvement in mean Dice score over both the original and fine-tuned models across two of the three evaluation splits. However, in the case of training on KISPI-IRTK and testing on KISPI-MIAL we notice a strong decrease in the performance of fine-tuned model (-2.5 DSC), showing the risk of overfitting to a particular training split when fine-tuning the model on real images. Nevertheless, by averaging the weights of the FetalSynthSeg and the fine-tuned model, we can recover from the performance dropped, and improve performance for other splits.
4.3.2 Target domain fine-tuning and weight averaging improve generalization to other domains
This last experiment investigated how using a few annotations from real images of a target domain impacts the generalization of FetalSynthSeg. We evaluated the fine-tuned model both on the target domain used for fine-tuning, and also on other unseed before domains. In Figure 11, we observe that weight averaging improves target domain performance compared to vanilla fine-tuning ($\alpha=1$), but also improves generalization to other domains, which does not occur with vanilla fine-tuning. Averaging weights with $\alpha=0.5$ or $0.8$ yields both a large target domain improvement (up to +5 in DSC with 3 images) which then translates into an improvement up to +1 DSC in other domains. Vanilla fine-tuning only improved performance in the target domain setting, while the performance in other domains remained unchanged, or even declined due to overfitting.
5 Discussion
This paper investigated how to best use synthetic data to train robust models that generalize out of domain. Building on SynthSeg (Billot et al., 2021, 2023a), our experiments explored the role of data generation, intensity clustering and real image fine-tuning in the context of single-source domain generalization using synthetic data. The challenges faced in fetal brain MRI, with datasets of limited size with large domain shifts (due to different acquisition methods and SR algorithms) are ubiquitous in medical imaging. Our rigorous evaluation of clinically relevant datasets, with both normal and pathological subjects allowed us to make observations that we believe will be of great help to practitioners looking to adapt SynthSeg-based methods to their data on new organs and modalities. In the discussion, we now revisit the main observations that we’ve drawn from our results.
(i) Intensity clustering is essential for a strong performance of domain randomization. In our experiments, we observed a large boost due to intensity clustering (around +5 in DSC), which was key for SynthSeg-based methods to outperform a model trained using real data, even if synthetically trained models remained more robust when confronted to out-of-domain volumes acquired at 0.55 T and reconstructed using unseen SR algorithms. This has initially been used in the cardiac context in (Billot et al., 2023b), and our results as well as others in the literature (Valabregue et al., 2024) confirm its crucial importance when few label classes are available.
(ii) Meta-classes can further improve the performance of the domain randomization method. In our experiments, meta-classes allowed FetalSynthSeg to get a small edge over SynthSeg ($\circ 1$ DSC point). They illustrate a good trade-off between pure domain randomization and physics-based constraints that can be imposed. If different classes contain the same type of tissues (e.g., CSF is found in both tissues around the brain and ventricles of the brain), their intensity distributions are expected to be similar within the same domain. Simulating different intensities for these tissues can then likely lead to the waste of model capacity.
(iii) Physics-based intensity generation methods do not improve performance compared to Gaussian mixture models When considering data generation procedures, it was clear that using FaBiAN (Lajous et al., 2022) with realistic physical parameters did not allow to achieve strong generalization performance. However, randFaBiAN, with randomized parameters, reached a much better performance of 68.4 DSC on average. This was still below SynthSeg (72.2 DSC), which benefited from a larger randomization by sampling intensities directly from a GMM. Indeed, it is easier to obtain large tissue intensity variations by directly sampling their values from a GMM than by randomizing tissue properties like T1 and T2 values in FaBiAN.
(iv) Data quality greatly impacts the generalization ability of synthetic methods. Models trained on KISPI-IRTK data consistently outperformed those trained on KISPI-MIAL data, which had worse data quality (Figure 2). This observation aligns with the better reported performance obtained with the BOUNTI method on their testing split, caused by the higher quality annotations (Uus et al., 2023), and the failure of SynthSeg-based approaches on poor quality data (Billot et al., 2023b; Valabregue et al., 2024).
By using intensity clustering, meta-classes, GMMs and good data quality, we were able to achieve a strong performance with FetalSynthSeg when dealing with large domain shifts: the method was able to generalize to data acquired on a low-field 0.55-T scanner, and reconstructed with an algorithm unseen during training (NeSVoR (Xu et al., 2023)). These findings show the limitation of training a model with only real, intensity-weighted images, where the real data model performed large errors that were avoided by FetalSynthSeg.
(v) Fine-tuning the model on a few segmentations from another domain boosts the performance of the model in-domain and also out-of-domain. Fine-tuning experiments showed the potential of fine-tuning the model on a few pairs of (image, segmentation) for a new domain (as demonstrated in other work (Lhermitte et al., 2024)), and using weight-averaging to improve the performance both on the new domain and for other domains. Techniques based on weight-averaging have seen great successes in computer vision and for large language models (Wortsman et al., 2022a; Rame et al., 2022; Ramé et al., 2023; Yadav et al., 2024), and could be promising ways to fine-tune models like SynthSeg and FetalSynthSeg in a community-driven manner, by only responsibly sharing the weights of fine-tuned models across multiple centers, without having to share any data. Indeed, contrarily to common fine-tuning, our results show that weight-averaging can bypass the trade-off between in-domain and out-of-domain performance: increasing in-domain performance with weight-averaging does not have to lead to a loss in out-of-domain performance.
Limitations and future work. While thorough, our experiments are limited to fetal brain MRI segmentation. Although we believe that it is an excellent test bed with datasets of limited size and strong domain shifts, it could be possible that not all of the observations will generalize to other settings. We have attempted to limit this issue with rigorous cross-validation and using a held-out pure testing set. As done by Billot et al. (2023a), we did not investigate the influence of model architectures on the generalization of the model. Although we expect these observations to hold independently of the model architecture chosen, there could be unexpected, more subtle dependencies. Beyond understanding more clearly how to train synthetic data model, future work will aim at building a stronger version of FetalSynthSeg, by leveraging a larger amount of data and multi-stage architectures like in (Billot et al., 2023b). This will be a first step towards developing robust segmentation tools that could be widely used across scanners, magnetic field strengths, and MR contrasts. Indeed, while our validation was limited to T2-weighted images, domain-randomized methods have shown to be robust to T1, T2, and other contrasts. Such robust, multi-contrast segmentation tools will be key to unlock more quantitative analysis tools in fetal neuroimaging and bridge the gap with neonatal imaging.
Declaration of competing interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data availability
The KISPI dataset is publicly available on its corresponding Synapse page (Payette and Jakab, 2021a). However, the KCL and CHUV data used for evaluation in this study are part of the private testing set from the FeTA challenge and are not publicly accessible.
CRediT author statement
Vladyslav Zalevskyi: Conceptualization, Investigation, Methodology, Software, Visualization, Writing – Original Draft Preparation. Thomas Sanchez: Conceptualization, Software, Visualization, Writing – Review & Editing, Supervision. Margaux Roulet: Software, Methodology, Writing – Original Draft Preparation. Hélène Lajous: Software, Methodology, Writing – Review & Editing. Jordina Aviles Verddera: Investigation, Data Curation. Jana Hutter: Investigation, Data Curation, Funding Acquisition. Hamza Kebiri: Writing – Review & Editing, Supervision. Meritxell Bach Cuadra: Funding Acquisition, Supervision, Writing – Review & Editing.
Acknowledgments
This research was funded by the Swiss National Science Foundation (182602 and 215641), ERA-NET Neuron MULTI-FACT project (SNSF 31NE30 203977), UKRI FLF (MR/T018119/1) and DFG Heisenberg funding (502024488); we acknowledge the Leenaards and Jeantet Foundations as well as CIBM Center for Biomedical Imaging, a Swiss research center of excellence founded and supported by CHUV, UNIL, EPFL, UNIGE and HUG.
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work, the authors used Grammarly and ChatGPT (GPT 4o) to assist with spell checking, grammar refinement, and language clarity improvements. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
Supplementary material
Experiment S1 – Ablation on data augmentations
In this additional experiment, we compared the SynthSeg-based augmentations to a common set of augmentations from MONAI (referred to as simple aug.). We re-trained the models using the following augmentations, each applied with a random probability of 0.5: random affine deformation (rotation=0.2, scale=0.1, translate=30, sheer=0.1), random contrast change (Gamma transformation) ($\gamma\in[0.5,1.5]$), random Gaussian noise ($\mu=0,\sigma=0.1$), random blurring ($\sigma\in[0.5,1.5]$), and scaling to the 0-1 range. Compared to these simple aug., SynthSeg-based augmentation has three additional components: a random non-linear deformation, random bias field, and a random resampling simulating an acquisition at a different resolution.
The results in Table 5 show a different picture than the results in the main paper. When comparing models trained using simple augmentations, the model trained using real data outperforms FetalSynthSeg (69.9 vs 66.9 DSC). Moving from simple augmentation increases the performance of all methods considered: FaBiAN (+2.9), randFaBiAN (+3.9), FetalSynthSeg (+8.0), Real Data (+1.9). This illustrates clearly how these data augmentation strategies are instrumental to the performance of FetalSynthSeg. They are also strong contenders for a model trained on real data and could be used in the future. A detailed ablation study of augmentations would be an interesting future step.
Table S1: Mean DSC for models trained on different sources of data. The variance is computed over all testing subjects within a split. The best performing method is shown in boldface in each column, and the second best is underlined.
{tblr}
[…]
randFaBiAN 74.1±3.1 69.0±6.4 63.8±10.3 54.0±9.3 62.0±13.8 62.8±13.0 64.3±12.1
FetalSynthSeg 72.9±4.070.3±6.271.4±9.368.3±8.258.7±15.560.0±16.1 66.9±11.8
[…]
randFaBiAN 79.1±2.3 78.2±2.9 55.1±12.7 68.9±7.8 60.7±7.3 68.1±14.7 68.3±13.8
SynthSeg 75.9±3.973.7±3.970.9±9.274.8±7.860.5±15.763.4±16.872.2±13.0
FetalSynthSeg 80.7±2.076.9±3.3 79.2±9.076.8±6.967.5±16.068.5±15.6 74.9±11.5
[…]
[r=3]c — Real — Real Data (Simple aug.) 76.5±3.275.2±3.569.6±13.667.7±12.967.2±17.663.4±17.0 69.9±12.6
Real Data (Synth. aug.)77.2±4.0 78.5±3.3 70.6±13.9 71.9±11.568.0±19.3 64.3±19.1 71.8±13.9
[…]
Experiment S2 – Comparing FaBiAN vs FetalSynthSeg on the same number of images
To ensure a fair comparison between FaBiAN and FetalSynthSeg we equalize the computational budget used to generate images for these two approaches. We generate offline 6000 synthetic images per split, as in our FaBiAN experiments, and use them to train a FetalSynthSeg-6k model. Table S2 shows that although this model achieves slightly lower results than FetalSynthSeg, it follows similar trends in performance across splits and still outperforms even the randFaBiAN approach overall (average Dice of FetalSynthSeg-6k 72.2±8.8 is vs average Dice of randFaBiAN of 68.3±8.5 across all splits).
Table S2: Dice scores comparing FetalSynthSeg model trained on the same amount of synthetic images as FaBiAN models.
{tblr}
row1 = c, row2 = c, cell12 = c=2, cell14 = c=2, cell16 = c=2, cell18 = r=2, hline1,7 = -0.08em, hline2 = 1-7, hline3 = -, Testing Split CHUV-mial KISPI-irtk KISPI-mial Global
Training Split KISPI-irtk KISPI-mial CHUV-mial KISPI-mial CHUV-mial KISPI-irtk
FaBiAN 74.2± 4.2 73.1± 5.7 53.6± 13.3 56.6± 12.9 60.6± 17.1 61.5± 20.1 63.3± 15.5
randFaBiAN 79.1± 2.3 78.2± 2.9 55.1± 12.7 68.9± 7.8 60.7± 7.3 68.1± 14.7 68.3± 13.8
FetalSynthSeg-6k 79.0± 2.6 76.0± 3.6 78.3± 9.2 75.1± 6.6 66.5± 15.9 67.7± 15.4 72.2± 13.0
FetalSynthSeg 80.7± 2.0 76.9± 3.3 79.2± 9.0 76.8± 6.9 67.5± 16.0 68.5± 15.6 74.9± 11.5
Additional qualitative results
We present additional qualitative results of the evaluated methods on several subjects across various domains. Figure S1 illustrates how the performance of models trained on the CHUV-MIAL dataset varies when tested on subjects from the KISPI-IRTK dataset across a wide range of gestational ages. Notably, as fetal age increases, cortical gray matter segmentation becomes more challenging due to the increasingly complex morphology.
Figure S1: Segmentation for subjects from the KISPI-IRTK dataset, showing the large age variability found across subjects.
Figure S2 shows the domain generalization capabilities of the tested models trained on different splits but evaluated on the two subjects from the same split. We can observe how the quality of the training data influences the models’ performance.
Figure S2: Generalization on two subjects from the CHUV-mial split. The top panel contains segmentation models trained using the KISPI-mial split, and the bottom panel contains segmentation models trained using the KISPI-IRTK split. This illustrates how models trained on different data generalize differently.
Additional quantitative results
We present additional quantitative results related to the conducted experiments, reporting the Dice scores per tissue for all of the tested models in Tables S3 as well as the average 95th-percentile Hausdorff distance score in Table S4.
Table S3: Mean and standard deviation of Dice score of the explored models for different testing and training splits. LV - lateral ventricles, CBM - cerebellum, SGM - sub-cortical gray matter, BS - brainstem, Mean DSC - value averaged across all segmentation labels.
{tblr}
row1 = c, cell21 = r=10c, cell22 = r=5c, cell72 = r=5c, cell121 = r=10c, cell122 = r=5c, cell172 = r=5c, cell221 = r=10c, cell222 = r=5c, cell272 = r=5c, hline1,32 = -0.08em, hline2,12,22 = -, hline7,17,27 = 2-11, Testing
 Splits
Training
 Split
Experiment Mean DSC CSF GM WM LV CBM SGM BS
CHUV-mial
KISPI-irtk FetalSynthSeg 80.7± 2.0 79.3± 2.7 70.7± 3.6 86.4± 2.3 78.7± 5.4 88.5± 2.3 79.6± 5.4 81.9± 2.7
Real data 77.2± 4.0 72.3± 7.1 68.4± 4.7 86.0± 2.8 75.8± 7.9 84.3± 9.2 76.4± 5.9 77.5± 4.7
FaBiAN 74.2± 4.2 75.3± 4.2 64.2± 4.8 81.3± 4.4 71.9± 9.9 85.3± 3.6 62.8± 14.2 78.7± 3.3
randFaBiAN 79.1± 2.3 77.0± 3.3 69.7± 3.6 85.9± 2.3 75.7± 5.3 87.4± 3.1 78.5± 8.4 79.5± 3.9
SynthSeg 75.9± 3.9 73.7± 4.5 64.1± 6.3 83.1± 4.4 71.9± 8.5 85.8± 2.9 74.0± 10.6 78.8± 3.9
KISPI-mial FetalSynthSeg 76.9± 3.3 73.1± 8.1 60.5± 4.8 84.9± 3.0 77.3± 6.3 86.2± 4.3 83.7± 4.6 72.8± 3.9
Real data 78.5± 3.3 79.0± 5.5 65.6± 4.4 85.8± 3.2 77.1± 7.3 86.5± 6.1 83.9± 3.8 71.4± 5.1
FaBiAN 73.1± 5.7 73.3± 4.9 61.9± 5.9 81.1± 4.8 66.3± 10.6 81.0± 14.9 76.3± 7.3 71.5± 7.1
randFaBiAN 78.2± 2.9 77.8± 3.9 64.1± 5.2 84.4± 3.2 75.5± 5.4 87.5± 4.6 83.0± 5.0 75.3± 5.0
SynthSeg 73.7± 4.4 67.9± 12.0 55.1± 8.7 81.3± 4.2 72.1± 7.2 84.0± 3.0 80.9± 5.2 74.6± 3.9
KISPI-irtk CHUV-mial FetalSynthSeg 79.2± 9.0 77.6± 10.9 69.4± 12.6 85.0± 12.9 79.7± 5.7 87.5± 9.1 74.8± 14.4 80.3± 6.6
Real data 70.6± 13.9 58.6± 18.4 60.9± 15.9 84.8± 9.8 72.8± 11.7 74.3± 26.3 74.5± 11.0 68.1± 19.9
FaBiAN 53.6± 13.3 51.4± 18.0 40.9± 11.9 74.9± 10.7 36.0± 13.8 45.9± 27.0 66.9± 12.1 59.0± 19.3
randFaBiAN 55.1± 12.7 54.2± 17.2 41.5± 11.5 76.9± 10.5 37.2± 13.7 51.8± 28.0 66.5± 11.7 57.8± 18.5
SynthSeg 70.9± 9.2 69.9± 11.6 58.6± 9.0 79.2± 11.8 66.4± 7.0 80.2± 11.3 68.8± 13.7 73.2± 9.2
KISPI-mial FetalSynthSeg 76.8± 6.9 78.7± 9.2 68.5± 9.1 86.5± 10.9 80.8± 4.7 85.1± 9.1 68.2± 12.3 69.9± 6.8
Real data 71.9± 11.5 72.3± 19.5 65.3± 13.4 85.5± 7.6 71.6± 9.0 80.3± 20.2 64.8± 13.3 63.2± 14.5
FaBiAN 56.6± 12.9 58.0± 17.6 39.3± 13.0 74.3± 9.5 50.6± 15.0 64.7± 27.7 58.7± 13.9 50.6± 18.7
randFaBiAN 68.9± 7.8 60.8± 16.0 52.8± 11.1 82.2± 6.3 68.1± 9.7 80.1± 10.9 68.0± 12.3 70.2± 7.9
SynthSeg 74.8± 7.8 74.4± 12.7 65.7± 10.8 85.4± 11.0 71.6± 7.3 86.3± 7.5 68.5± 13.1 72.0± 8.0
KISPI-mial
CHUV-mial FetalSynthSeg 67.5± 16.0 59.5± 28.3 48.7± 19.4 74.3± 17.3 78.6± 12.3 67.2± 29.7 79.3± 9.8 64.9± 16.0
Real data 68.0± 19.3 62.1± 32.4 58.3± 15.7 81.8± 12.3 77.0± 14.5 64.6± 37.3 78.8± 9.9 53.4± 29.8
FaBiAN 60.6± 17.1 52.9± 29.1 48.7± 15.2 77.6± 12.3 63.2± 17.3 53.7± 33.1 73.4± 14.5 54.8± 25.8
randFaBiAN 60.7± 16.3 52.3± 28.6 49.6± 14.5 78.2± 12.4 63.7± 18.0 52.8± 32.5 73.8± 14.3 54.6± 22.4
SynthSeg 60.5± 15.7 53.7± 26.0 43.8± 14.1 70.0± 15.4 65.9± 12.1 59.9± 28.5 71.1± 15.5 59.0± 17.0
KISPI-irtk FetalSynthSeg 68.5± 15.6 63.2± 25.9 57.2± 16.8 82.1± 13.5 79.4± 12.2 73.2± 24.9 63.6± 18.4 60.8± 17.0
Real data 64.3± 19.1 59.8± 29.0 55.0± 17.9 79.1± 15.5 72.6± 12.9 64.4± 33.8 62.5± 16.3 57.1± 21.6
FaBiAN 61.5± 20.1 57.5± 31.5 53.1± 19.0 80.6± 11.8 72.3± 16.1 59.1± 35.1 52.8± 20.9 54.9± 23.4
randFaBiAN 68.1± 14.7 62.5± 27.3 59.4± 12.4 83.8± 12.2 80.4± 11.0 70.6± 25.0 62.3± 21.2 57.4± 15.1
SynthSeg 63.4± 16.8 57.3± 26.8 50.2± 17.6 77.4± 16.0 77.7± 13.4 70.4± 22.0 54.7± 23.9 56.2± 18.2
Table S4: Mean and standard deviation of 95-th percentile Hasudorff distance for the explored models on different testing and training splits. LV - lateral ventricles, CBM - cerebellum, SGM - sub-cortical gray matter, BS - brainstem, mHD95 - average value across all labels.
{tblr}
row1 = c, cell21 = r=10c, cell22 = r=5c, cell72 = r=5c, cell121 = r=10c, cell122 = r=5c, cell172 = r=5c, cell221 = r=10c, cell222 = r=5c, cell272 = r=5c, hline1,32 = -0.08em, hline2,12,22 = -, hline7,17,27 = 2-11, Testing
Splits
Training
Split
Experiment mHD95 CSF GM WM Ventricles Cerebellum Deep_GM Brainstem
CHUV - mial
KISPI-irtk FetalSynthSeg 2.0± 0.5 1.7± 0.3 1.6± 0.3 2.5± 3.3 2.0± 0.9 1.5± 0.2 2.8± 0.6 2.1± 0.6
Real data 2.9± 1.2 2.6± 0.8 1.9± 0.5 2.0± 0.3 4.2± 4.4 2.9± 4.7 4.0± 2.8 2.8± 1.1
FaBiAN 3.6± 1.3 2.6± 0.9 2.2± 0.4 3.6± 0.8 5.3± 4.1 3.5± 6.0 5.0± 2.3 2.9± 1.0
randFaBiAN 2.1± 0.3 1.8± 0.3 1.6± 0.3 2.1± 0.3 2.1± 0.7 1.6± 0.4 3.0± 0.8 2.6± 1.0
SynthSeg 2.7± 0.6 2.0± 0.4 2.1± 1.9 2.5± 0.3 3.7± 1.8 2.3± 0.8 3.3± 0.8 2.9± 1.4
KISPI-mial FetalSynthSeg 3.1± 0.7 2.7± 1.7 2.5± 0.5 2.2± 0.4 2.6± 1.7 2.9± 2.9 2.6± 0.5 6.3± 0.9
Real data 3.0± 0.7 2.2± 1.6 2.3± 0.4 2.2± 0.5 3.1± 3.0 1.9± 0.8 2.9± 0.6 6.6± 1.1
FaBiAN 4.1± 1.4 3.1± 1.9 2.8± 0.6 4.0± 1.2 4.6± 3.4 2.8± 1.7 5.7± 5.2 5.4± 1.3
randFaBiAN 2.7± 0.6 1.9± 0.6 2.2± 0.5 2.4± 0.5 2.1± 0.6 2.5± 2.2 2.8± 0.7 5.0± 1.2
SynthSeg 3.4± 0.6 3.5± 2.9 2.4± 0.4 2.9± 0.3 4.6± 1.9 3.1± 1.2 3.2± 0.6 3.9± 0.9
KISPI - irtk
CHUV-mial FetalSynthSeg 2.5± 1.6 2.4± 2.2 1.6± 0.8 2.2± 1.4 1.9± 1.0 1.5± 0.9 4.3± 4.0 3.4± 3.6
Real data 9.8± 5.1 9.0± 2.3 7.6± 4.3 5.7± 4.2 6.3± 6.0 16.1± 16.1 9.5± 9.6 14.1± 9.7
FaBiAN 14.8± 5.3 9.2± 2.1 10.3± 2.2 9.6± 2.7 14.2± 4.3 28.7± 16.6 15.1± 8.7 16.3± 11.1
randFaBiAN 14.8± 5.1 9.0± 2.0 10.7± 2.3 10.0± 3.2 13.2± 5.5 24.4± 16.5 17.1± 10.1 18.9± 11.3
SynthSeg 1.5± 0.5 1.3± 1.0 1.2± 0.3 1.4± 0.4 1.4± 0.6 1.3± 0.8 2.0± 0.8 1.7± 0.7
KISPI-mial FetalSynthSeg 3.4± 1.7 2.4± 1.5 1.7± 0.8 2.1± 0.7 3.1± 3.4 3.4± 6.5 5.5± 2.9 5.9± 1.4
Real data 5.9± 4.0 6.3± 4.0 4.1± 4.8 5.4± 6.8 4.3± 4.8 6.3± 10.8 6.8± 4.2 7.8± 4.7
FaBiAN 14.0± 5.0 9.0± 2.1 10.1± 2.7 11.1± 3.4 8.9± 6.3 24.9± 18.4 18.6± 11.0 15.4± 9.1
randFaBiAN 5.5± 2.6 6.8± 2.6 3.9± 1.9 3.8± 2.3 5.3± 5.6 6.7± 10.0 6.3± 5.6 5.8± 3.2
SynthSeg 3.5± 1.4 2.9± 2.0 1.7± 0.8 2.3± 0.7 3.8± 3.2 3.7± 6.3 5.3± 2.0 4.4± 1.3
KISPI - mial
CHUV-mial FetalSynthSeg 4.1± 2.4 4.8± 4.5 3.2± 1.9 3.4± 1.8 2.4± 2.3 4.5± 3.9 3.8± 3.3 6.4± 3.3
Real data 5.1± 3.4 5.0± 4.9 3.1± 1.7 2.8± 1.3 4.1± 4.1 7.7± 9.8 4.1± 3.0 8.6± 6.3
FaBiAN 8.2± 4.3 7.3± 4.9 6.5± 2.4 4.1± 1.4 8.9± 4.4 11.2± 12.4 7.3± 7.7 12.3± 8.7
randFaBiAN 7.6± 4.2 7.0± 4.8 6.0± 1.8 3.9± 1.4 8.0± 4.4 9.9± 10.6 7.8± 8.1 10.5± 8.1
SynthSeg 2.6± 1.6 3.1± 3.2 2.2± 2.2 2.6± 3.0 1.7± 1.0 2.5± 1.7 2.7± 3.4 3.6± 4.3
KISPI-irtk FetalSynthSeg 3.9± 2.0 3.9± 3.6 2.4± 1.5 2.6± 1.3 2.3± 2.0 3.6± 3.1 5.6± 2.6 6.7± 2.7
Real data 4.7± 2.7 5.2± 4.6 2.8± 1.7 2.7± 1.4 4.7± 3.9 5.9± 6.5 5.8± 2.6 5.9± 1.8
FaBiAN 6.5± 3.9 6.5± 4.9 4.7± 3.3 4.4± 2.4 6.4± 5.4 8.3± 7.9 7.0± 5.5 8.4± 6.0
randFaBiAN 4.1± 2.0 4.4± 4.3 2.8± 1.3 3.1± 2.8 2.2± 1.7 3.5± 2.3 5.7± 2.8 7.0± 2.6
SynthSeg 5.2± 2.8 4.9± 4.2 3.6± 2.0 3.7± 3.0 4.4± 6.5 5.1± 8.1 6.7± 3.9 7.9± 2.9
Generation parameters
In table S5 we report parameters used to create synthetic images with the the FaBiAN generator.
[…]
Echo spacing (ms) 6.12
Echo train length 150
Excitation flip angle (°) 90
Refocusing pulse flip angle (°) [150,180]
[…]
Slice gap (mm) 0
Number of slices 112
Phase oversampling (%) 80
Shift of field-of-view (mm) 0
[…]
Base resolution (voxels) 150
Phase resolution (%) 70
Reconstruction matrix 150
Zero-interpolation filling 1
[…]
Acceleration factor 2
[…]
Standard deviation <0.01
[…]
Table S5: FaBiAN Simulation Parameters for HASTE Imaging